Write the following rational numbers in ascending order: -3/7, -3/2, -3/4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

To convert the given rational numbers into like fraction we have to find LCM, LCM of 7, 2, and 4 is 28
Now,
(-3/7)= [(-3 × 4)/ (7 × 4)] = (-12/28)
(-3/2)= [(-3 × 14)/ (2 × 14)] = (-42/28)
(-3/4)= [(-3 × 7)/ (4 × 7)] = (-21/28)
Clearly,
(-42/28) < (-21/28) < (-12/28)
Hence,
(-3/2) < (-3/4) < (-3/7)

​​Thanks